Find the equation of the sphere with the given conditions:


(a) Center (2,4,3), tangent plane 2x-3y+4z-5=0


(b) endpoints of the diameter are (-5,3,2) and (-1,-3,2)


1
Expert's answer
2021-02-24T07:49:00-0500

(a) Let us find the equation of the sphere with the center "(2,4,3)" and tangent plane "2x-3y+4z-5=0". The radius of the sphere is equal to the distance from the center to the tangent plane: "R=\frac{|2\cdot 2-3\cdot 4+4\cdot 3-5|}{\sqrt{2^2+(-3)^2+4^2}}=\frac{1}{\sqrt{29}}." Consequently, the equation of the sphere is the following:


"(x-2)^2+(y-4)^2+(z-3)^2=\frac{1}{29}"


(b) Let us find the equation of the sphere in the case that the endpoints of the diameter are "(-5,3,2)" and "(-1,-3,2)" . It follows that the center of the sphere is "O(a,b,c)", where "a=\frac{-5-1}{2}=-3,\ b=\frac{3-3}{2}=0,\ c=\frac{2+2}{2}=2." The radius is equal the distance from the center to "(-1,-3,2):" "R=\sqrt{(-1+3)^2+(-3-0)^2+(2-2)^2}=\sqrt{13}". Therefore, the equation of the sphere is the following:


"(x+3)^2+y^2+(z-2)^2=13"